# Attention Needs to Focus: A Unified Perspective on Attention Allocation

### v1

@@ 1 Introduction
The remarkable success of Large Language Models (LLMs) across a diverse range of tasks, from text generation (gpt2) to complex reasoning (cot), hinges on the Transformer architecture’s core innovation: the self-attention mechanism. By computing pairwise similarity scores between queries and keys, self-attention enables a model to dynamically construct context-aware representations. The underlying principle is to assign greater weight to semantically relevant tokens and lesser weight to irrelevant ones, thereby enabling the model to capture meaningful dependencies across tokens in the sequence.
However, empirical evidence shows that the standard attention mechanism often deviates from this ideal, giving rise to two widely observed phenomena: representational collapse (squashing) and attention sink (streamingllm). Representational collapse occurs in relatively long-context scenarios, where the attention mechanism tries to aggregate information from too many tokens, leading to over-squashing and ultimately indistinguishable final representations. Conversely, attention sink describes that the initial tokens disproportionately attract high attention weights and serve as anchors to stabilize the attention distribution even though they carry little semantic significance (whensinkemerge). Both phenomena degrade model performance and hinder efficient deployment, suggesting a deeper structural deficiency in current attention design.
Existing works address these problems separately. For the collapse issue, common approaches increase model dimensionality or numerical precision, which alleviates representation loss but at the cost of significantly higher training and inference overhead. For the sink phenomenon, prevailing solutions adapt to this by explicitly preserving a sink token (duoattention) or sink bias (gptoss), which reduces model flexibility, or by removing normalization, which often comes at the expense of performance stability. Although such approaches mitigate individual issues, they are inherently fragmented and fail to handle both problems within a unified framework.
In this paper, we provide a unified perspective on the above problems: the core limitation of self-attention lies in its improper attention allocation, manifesting in two opposite extremes (Figure 1). Attention Overload occurs when dense contexts spread attention too broadly across tokens, causing semantic features to be averaged away and resulting in indistinguishable representations (Efficient). Attention Underload when a token has little relevance to the previous context, yet softmax normalization requires the attention weights to sum to one, leading to the sink phenomenon. Together, these two failure modes highlight the need for an attention mechanism that allocates focus more selectively, amplifying the attention weight of informative tokens while ignoring irrelevant ones.
@@ 2.1 Transformer Architecture
A Transformer (attention) is composed of $L$ stacked layers, each consisting of a multi-head attention (MHA) module and a feed-forward network (FFN), both equipped with residual connections. Given an input sequence $\bm{X}^{l}=[\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{n}]\in\mathbb{R}^{n\times d}$ at layer $l$, where $n$ is the sequence length and $d$ the hidden dimension, the layer update can be written as:
@@ 2.2 Position Encoding
Since self-attention is permutation-invariant, positional information must be incorporated into the model. Transformer (attention) introduced sinusoidal functions to create fixed positional encodings, which often referred to as absolute position encodings (APE), are defined as $\bm{P}_{(i,2t)}=\sin(i/10000^{2t/d})$ and $\bm{P}_{(i,2t+1)}=\cos(i/10000^{2t/d})$. where $i$ is the token position, $t$ is the dimension and $d$ is the hidden dimension size. Subsequent works like GPT (gpt2) adopted learnable position embeddings provided by model parameters. In both variants, position embeddings $\bm{P}\in\mathbb{R}^{n\times d}$ are added to token embeddings: $\text{TokenEmbed}(\bm{X})+\bm{P}$ and input to the first Transformer layer. However, APEs have poor length generalization beyond training sequences.
To address the limitations of APE, relative position encoding (RPE) was introduced to model the pairwise distance between tokens. Instead of adding embeddings, RPEs typically modify the attention mechanism itself, especially the attention scores. Influential methods include Transformer-XL (transformer-xl), which uses a granular attention formula that disentangles content and relative position, and ALiBi (alibi), which simply adds a static bias (e.g., $m\cdot|i-j|$) to the attention scores. More recently, rotary position embedding (RoPE) (rope) has become adopted in LLMs. It encodes relative position by applying rotational transformations, concisely represented as $\mathbf{q}^{\prime}_{i}=\bm{R}_{i}\mathbf{q}_{i},\;\mathbf{k}^{\prime}_{i}=\bm{R}_{i}\mathbf{k}_{i}$, where $\bm{R}_{i}$ is a block-diagonal rotation matrix, allowing attention scores to reflect relative positions. More related works are detailed in Appendix B.1
@@ 3.1 The Nature of Attention Sink
Attention sink refers to the phenomenon that initial tokens attract a large amount of attention during inference (streamingllm; lm-infinite). Related discussions are provided in Appendix B.2.
Following whensinkemerge, we probe the location of the attention sink by inserting a fixed [Mask] token at a specific position during pre-training, denoted as “Mask@k”, where k indicates its position in the input. As shown in Figure 2(a), when the fixed token is placed near the beginning (e.g., Mask@2), it attracts attention from the first position. However, when inserted beyond approximately 16 tokens from the start (e.g., Mask@16), the first token regains the sink behavior.
Meanwhile, varience pointed out that the variance of hidden states in Transformers contains latent positional information even without positional embeddings. This is consistent with the comparison of Figure 2(b) and 2(c). When attention sink exists, the variance of the sink token’s hidden states is much greater than that of other tokens. Similarly, the variance of the sink token’s value in the attention calculation is much smaller than that of other tokens, as detailed in Appendix G.
@@ 3.2 The Role of Position Encoding
With APE (e.g., GPT-2), the sink phenomenon persists even under repeated inputs (Figure 3b.1). In contrast, LLMs adopting RPE—such as Llama2 (llama2) with RoPE and Bloom (bloom) with ALiBi—lose their sink behavior entirely. Their attention distributions collapse into uniform patterns modulated by long-term decays of position encodings. This indicates that RoPE, ALiBi, and similar RPE methods fundamentally provide position-sensitive attention weights rather than embed positional information into hidden states that can propagate across layers.
@@ 4.2 Elastic-Softmax
where $\tau^{(h)}$ is a learnable head-specific offset, initialized to 1-1 and divided evenly across the $i$ attended tokens. ReLU (relu) ensures all weights remain non-negative.
Elastic-Softmax eliminates insignificant attention weights, resulting in sparse attention that mitigates attention sink and emphasizes informative tokens. For instance, if no specific token stands out as a focus of attention, attention weights are offset to zero when $\tau^{(h)}=-1$. Moreover, the formulation is fully compatible with FlashAttention (flash), as detailed in Appendix F.
@@ 5.1 Experiment Settings
Datasets. LLMs are trained on FineWeb-Edu (fineweb-edu) 10BT and 100BT; 15B/30B are taken sequentially from shuffled 100BT, and ablations use full 10BT (see Appendix C.1 for details).
Baselines. We benchmark our method against: (i) Transformer variants: Transformer++ (llama2) and Sigmoid Attention (sigmoid); (ii) Recurrent architectures: RetNet (retnet), Mamba2 (mamba2), TTT (ttt), Gated DeltaNet (gateddeltanet), and Titans (titans). (iii) Streaming inference: StreamingLLM (streamingllm).
@@ 5.3 What Long-term Decay Does Attention Really Need?
Inhibition of Return (IOR). The upper layers display a non-monotonic bias characterized by three unique patterns: (i) positive bias for immediate neighbors, followed by a sharp descent to negative values (1-10), (ii) gradual recovery (10-100) to zero, and (iii) stabilization near zero for all distant positions. This structure resembles inhibition-of-return (IOR) (ior) in human visual attention, where recently attended locations are temporarily suppressed to promote exploration. This spontaneous emergence indicates that resolving overload by suppressing redundant mid-range attention while preserving access to long-range dependencies benefits language modeling.
@@ B.1 Position Encoding
Position encoding is essential in Transformer models since self-attention itself is permutation-invariant. Early studies mainly adopted absolute position embeddings (APE). The original Transformer used fixed sinusoidal functions, and GPT-2 (gpt2) later replaced them with trainable vectors fused with token representations. Although these approaches worked well for sequences within the training length, they struggled to generalize once the context became longer.
Relative position encoding (RPE) offered a more flexible alternative. Transformer-XL (transformer-xl) introduced learnable biases based on token distance, which allowed recurrent processing between segments. ALiBi (alibi) simplified the idea by applying fixed linear biases that decay with distance, making it possible to extrapolate to much longer inputs. These methods shifted attention from absolute indices to relative distances, laying the groundwork for position encodings that are better suited to scaling Transformers toward long and variable contexts.
Rotary Position Embeddings (RoPE) (rope) marked another step forward by encoding relative positions as rotations applied to the query and key vectors. RoPE has since become standard in many LLMs. Building on this, several extensions have been proposed. For example, YaRN rescales the rotation angles to push LLaMA models to million-token contexts with limited retraining (yarn). LongRoPE2 searches for scaling factors through evolutionary methods to balance precision in both short and long ranges (longrope2). XPOS combines blockwise causal attention with midrange relative biases to improve interpolation stability (xpos). Other designs, such as Fourier Position Embedding (FoPE) (fope), emphasize frequency filtering to retain periodic structure and help models remain stable when context length grows.
@@ B.2 Attention Sink
Attention sink refers to the phenomenon that initial tokens attract a large amount of attention during inference (streamingllm; lm-infinite). Subsequent analysis links this to normalization, which forces LLMs to assign attention even when no informative value (whensinkemerge).
To mitigate potential issues associated with this phenomenon, various solutions have been developed. Most models try to adapt this pattern by leaving the first token as the sink token and always keeping it during inference (duoattention; minference). Similarly, GPT-OSS (gptoss) adds a learnable parameter to the denominator of the softmax function, which is equal to create a “hidden” sink token with constant attraction, thereby alleviating the sink issue without requiring a physical placeholder token. Sigmoid attention (sigmoid) can completely avoid attention sink while sacrificing model performance and training stability.
@@ B.3 Sparse and Efficient Attention
Another challenge comes from the quadratic cost of self-attention. Fixed-pattern sparsity, exemplified by Longformer (Longformer), adopts sliding windows with global tokens, which improves efficiency but reduces flexibility. Linformer (lineartransformer) takes a different route by approximating attention with low-rank projections, reducing the cost to linear time.
More recent work has looked at dynamic and normalization based mechanisms. Among them, StreamingLLM (streamingllm) introduced the notion of an “attention sink,” caching only a few initial tokens to enable million-token inference even for models trained on short windows. Sigmoid Attention (sigmoid) implemented a sigmoid variant of attention in a hardware-friendly way, achieving more than 15% faster inference without sacrificing accuracy. These approaches illustrate how relatively small design changes can make long-context inference practical.
A complementary direction is cache optimization. DuoAttention (duoattention) splits heads into retrieval and streaming roles, keeping full key–value caches only for retrieval to reduce memory. RetroAttention (retroattention) refreshes caches with new entries to improve coverage of early tokens. FlexPrefill (flexprefill) adapts the prefill stage by selecting query-aware sparse indices, balancing accuracy and efficiency during initialization. Together, these methods show how careful memory management can extend usable context length without prohibitive compute costs.
@@ C.1 Datasets
All experiments use subsets of FineWeb-Edu (fineweb-edu). We train on 10B-token (10BT) and 100B-token (100BT) pools; preprocessing applies a global shuffle (seed=42). The 15B and 30B budgets are obtained deterministically by taking the first 15B / 30B tokens from the start of the shuffled 100BT subset. Ablation runs use the full 10BT pool, which is shuffled once in preprocessing and then used with a fixed order during training. For all model sizes and runs the training data and its order are identical (the same shuffled are used), ensuring data-level consistency across experiments.
@@ C.2 Implementation details
All models and baselines are implemented on top of the FLA (fla) and Flame (flame) stack, which provides Triton-based efficient-attention kernels and standard LM training utilities. For sigmoid-attention experiments we replace the FlashAttention kernel with the authors’ Flash-Sigmoid implementation integrated into the same stack.
@@ C.4 Benchmarks
Wikitext (wikitext): A large linguistic corpus extracted from Wikipedia articles, containing over 100 million word tokens. It tests a model’s ability to predict the next word in a passage of text.
Lambada (lambada): The LAmBdA dataset tests a model’s capability of using broad discourse context to predict the last word of a passage extracted from books. It contains over 60,000 examples.
PIQA (PIQA): The Physical Interaction: Question Answering (PIQA) dataset tests commonsense reasoning about physical interactions between two entities. It contains 16,113 multiple choice questions generated from crowd-sourcing.
Hellaswag (Hellaswag): The HellaSwag dataset consists of 70,000 multiple choice questions about inferring what might happen next in a story. It requires commonsense reasoning to choose the most plausible ending.
WinoGrande (WinoGrande): The WinoGrande dataset tests coreference resolution and commonsense reasoning with 44,000 examples obtained from books and websites.
ARC (arc): The AI2 Reasoning Challenge (ARC) dataset contains 7,787 genuine grade-school level, multiple-choice science questions, grouped into an Easy Set (ARC-e) and a Challenge Set (ARC-c).
SIQA (siqa): The Social Interaction QA (SIQA) dataset contains 15,554 multiple choice questions that describe situations about people’s social interactions.
BoolQ (boolq): The Boolean Questions (BoolQ) dataset contains 15,942 English yes/no questions sampled from Google search queries to test a model’s ability to answer simple questions.

### v2

@@ 1 Introduction
The remarkable success of Large Language Models (LLMs) across a diverse range of tasks, from text generation (Radford et al., 2019) to complex reasoning (Wei et al., 2022), hinges on the Transformer architecture’s core innovation: the self-attention mechanism. By computing pairwise similarity scores between queries and keys, self-attention enables a model to dynamically construct context-aware representations. The underlying principle is to assign greater weight to semantically relevant tokens and lesser weight to irrelevant ones, thereby enabling the model to capture meaningful dependencies across tokens in the sequence.
However, empirical evidence shows that the standard attention mechanism often deviates from this ideal, giving rise to two widely observed phenomena: representational collapse (Barbero et al., 2024) and attention sink (Xiao et al., 2023). Representational collapse occurs in relatively long-context scenarios, where the attention mechanism tries to aggregate information from too many tokens, leading to over-squashing and ultimately indistinguishable final representations. Conversely, attention sink describes that the initial tokens disproportionately attract high attention weights and serve as anchors to stabilize the attention distribution even though they carry little semantic significance (Gu et al., 2024). Both phenomena degrade model performance and hinder efficient deployment, suggesting a deeper structural deficiency in current attention design.
Existing works address these problems separately. For the collapse issue, common approaches increase model dimensionality or numerical precision, which alleviates representation loss but at the cost of significantly higher training and inference overhead. For the sink phenomenon, prevailing solutions adapt to this by explicitly preserving a sink token (Xiao et al., 2025) or sink bias (OpenAI et al., 2025), which reduces model flexibility, or by removing normalization, which often comes at the expense of performance stability. Although such approaches mitigate individual issues, they are inherently fragmented and fail to handle both problems within a unified framework.
In this paper, we provide a unified perspective on the above problems: the core limitation of self-attention lies in its improper attention allocation, manifesting in two opposite extremes (Figure 1). Attention Overload occurs when dense contexts spread attention too broadly across tokens, causing semantic features to be averaged away and resulting in indistinguishable representations (Sun et al., 2025). Attention Underload when a token has little relevance to the previous context, yet softmax normalization requires the attention weights to sum to one, leading to the sink phenomenon. Together, these two failure modes highlight the need for an attention mechanism that allocates focus more selectively, amplifying the attention weight of informative tokens while ignoring irrelevant ones.
@@ 2.1 Transformer Architecture
A Transformer (Vaswani et al., 2017) is composed of $L$ stacked layers, each consisting of a multi-head attention (MHA) module and a feed-forward network (FFN), both equipped with residual connections. Given an input sequence $\bm{X}^{l}=[\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{n}]\in\mathbb{R}^{n\times d}$ at layer $l$, where $n$ is the sequence length and $d$ the hidden dimension, the layer update can be written as:
@@ 2.2 Position Encoding
Since self-attention is permutation-invariant, positional information must be incorporated into the model. Transformer (Vaswani et al., 2017) introduced sinusoidal functions to create fixed positional encodings, which often referred to as absolute position encodings (APE), are defined as $\bm{P}_{(i,2t)}=\sin(i/10000^{2t/d})$ and $\bm{P}_{(i,2t+1)}=\cos(i/10000^{2t/d})$. where $i$ is the token position, $t$ is the dimension and $d$ is the hidden dimension size. Subsequent works like GPT (Radford et al., 2019) adopted learnable position embeddings provided by model parameters. In both variants, position embeddings $\bm{P}\in\mathbb{R}^{n\times d}$ are added to token embeddings: $\text{TokenEmbed}(\bm{X})+\bm{P}$ and input to the first Transformer layer. However, APEs have poor length generalization beyond training sequences.
To address the limitations of APE, relative position encoding (RPE) was introduced to model the pairwise distance between tokens. Instead of adding embeddings, RPEs typically modify the attention mechanism itself, especially the attention scores. Influential methods include Transformer-XL (Dai et al., 2019), which uses a granular attention formula that disentangles content and relative position, and ALiBi (Press et al., 2022), which simply adds a static bias (e.g., $m\cdot|i-j|$) to the attention scores. More recently, rotary position embedding (RoPE) (Su et al., 2023) has become adopted in LLMs. It encodes relative position by applying rotational transformations, concisely represented as $\mathbf{q}^{\prime}_{i}=\bm{R}_{i}\mathbf{q}_{i},\;\mathbf{k}^{\prime}_{i}=\bm{R}_{i}\mathbf{k}_{i}$, where $\bm{R}_{i}$ is a block-diagonal rotation matrix, allowing attention scores to reflect relative positions. More related works are detailed in Appendix B.1
@@ 3.1 The Nature of Attention Sink
Attention sink refers to the phenomenon that initial tokens attract a large amount of attention during inference (Xiao et al., 2023; Han et al., 2024). Related discussions are provided in Appendix B.2.
Following Gu et al. (2024), we probe the location of the attention sink by inserting a fixed [Mask] token at a specific position during pre-training, denoted as “Mask@k”, where k indicates its position in the input. As shown in Figure 2(a), when the fixed token is placed near the beginning (e.g., Mask@2), it attracts attention from the first position. However, when inserted beyond approximately 16 tokens from the start (e.g., Mask@16), the first token regains the sink behavior.
Meanwhile, Chi et al. (2023) pointed out that the variance of hidden states in Transformers contains latent positional information even without positional embeddings. This is consistent with the comparison of Figure 2(b) and 2(c). When attention sink exists, the variance of the sink token’s hidden states is much greater than that of other tokens. Similarly, the variance of the sink token’s value in the attention calculation is much smaller than that of other tokens, as detailed in Appendix G.
@@ 3.2 The Role of Position Encoding
With APE (e.g., GPT-2), the sink phenomenon persists even under repeated inputs (Figure 3b.1). In contrast, LLMs adopting RPE—such as Llama2 (Touvron et al., 2023) with RoPE and Bloom (Scao et al., 2022) with ALiBi—lose their sink behavior entirely. Their attention distributions collapse into uniform patterns modulated by long-term decays of position encodings. This indicates that RoPE, ALiBi, and similar RPE methods fundamentally provide position-sensitive attention weights rather than embed positional information into hidden states that can propagate across layers.
@@ 4.2 Elastic-Softmax
where $\tau^{(h)}$ is a learnable head-specific offset, initialized to 1-1 and divided evenly across the $i$ attended tokens. ReLU (Agarap, 2018) ensures all weights remain non-negative.
Elastic-Softmax eliminates insignificant attention weights, resulting in sparse attention that mitigates attention sink and emphasizes informative tokens. For instance, if no specific token stands out as a focus of attention, attention weights are offset to zero when $\tau^{(h)}=-1$. Moreover, the formulation is fully compatible with FlashAttention (Dao et al., 2022), as detailed in Appendix F.
@@ 5.1 Experiment Settings
Datasets. LLMs are trained on FineWeb-Edu (Lozhkov et al., 2024) 10BT and 100BT; 15B/30B are taken sequentially from shuffled 100BT, and ablations use full 10BT (see Appendix C.1 for details).
Baselines. We benchmark our method against: (i) Transformer variants: Transformer++ (Touvron et al., 2023) and Sigmoid Attention (Ramapuram et al., 2025); (ii) Recurrent architectures: RetNet (Sun et al., 2023a), Mamba2 (Dao and Gu, 2024), TTT (Sun et al., 2024), Gated DeltaNet (Yang et al., 2024), and Titans (Behrouz et al., 2024). (iii) Streaming inference: StreamingLLM (Xiao et al., 2023).
@@ 5.3 What Long-term Decay Does Attention Really Need?
Inhibition of Return (IOR). The upper layers display a non-monotonic bias characterized by three unique patterns: (i) positive bias for immediate neighbors, followed by a sharp descent to negative values (1-10), (ii) gradual recovery (10-100) to zero, and (iii) stabilization near zero for all distant positions. This structure resembles inhibition-of-return (IOR) (Klein and Ivanoff, 2000) in human visual attention, where recently attended locations are temporarily suppressed to promote exploration. This spontaneous emergence indicates that resolving overload by suppressing redundant mid-range attention while preserving access to long-range dependencies benefits language modeling.
@@ B.1 Position Encoding
Position encoding is essential in Transformer models since self-attention itself is permutation-invariant. Early studies mainly adopted absolute position embeddings (APE). The original Transformer used fixed sinusoidal functions, and GPT-2 (Radford et al., 2019) later replaced them with trainable vectors fused with token representations. Although these approaches worked well for sequences within the training length, they struggled to generalize once the context became longer.
Relative position encoding (RPE) offered a more flexible alternative. Transformer-XL (Dai et al., 2019) introduced learnable biases based on token distance, which allowed recurrent processing between segments. ALiBi (Press et al., 2022) simplified the idea by applying fixed linear biases that decay with distance, making it possible to extrapolate to much longer inputs. These methods shifted attention from absolute indices to relative distances, laying the groundwork for position encodings that are better suited to scaling Transformers toward long and variable contexts.
Rotary Position Embeddings (RoPE) (Su et al., 2023) marked another step forward by encoding relative positions as rotations applied to the query and key vectors. RoPE has since become standard in many LLMs. Building on this, several extensions have been proposed. For example, YaRN rescales the rotation angles to push LLaMA models to million-token contexts with limited retraining (Peng et al., 2024). LongRoPE2 searches for scaling factors through evolutionary methods to balance precision in both short and long ranges (Shang et al., 2025). XPOS combines blockwise causal attention with midrange relative biases to improve interpolation stability (Sun et al., 2023b). Other designs, such as Fourier Position Embedding (FoPE) (Hua et al., 2024), emphasize frequency filtering to retain periodic structure and help models remain stable when context length grows.
@@ B.2 Attention Sink
Attention sink refers to the phenomenon that initial tokens attract a large amount of attention during inference (Xiao et al., 2023; Han et al., 2024). Subsequent analysis links this to normalization, which forces LLMs to assign attention even when no informative value (Gu et al., 2024).
To mitigate potential issues associated with this phenomenon, various solutions have been developed. Most models try to adapt this pattern by leaving the first token as the sink token and always keeping it during inference (Xiao et al., 2025; Jiang et al., 2024). Similarly, GPT-OSS (OpenAI et al., 2025) adds a learnable parameter to the denominator of the softmax function, which is equal to create a “hidden” sink token with constant attraction, thereby alleviating the sink issue without requiring a physical placeholder token. Sigmoid attention (Ramapuram et al., 2025) can completely avoid attention sink while sacrificing model performance and training stability.
@@ B.3 Sparse and Efficient Attention
Another challenge comes from the quadratic cost of self-attention. Fixed-pattern sparsity, exemplified by Longformer (Beltagy et al., 2020), adopts sliding windows with global tokens, which improves efficiency but reduces flexibility. Linformer (Katharopoulos et al., 2020) takes a different route by approximating attention with low-rank projections, reducing the cost to linear time.
More recent work has looked at dynamic and normalization based mechanisms. Among them, StreamingLLM (Xiao et al., 2023) introduced the notion of an “attention sink,” caching only a few initial tokens to enable million-token inference even for models trained on short windows. Sigmoid Attention (Ramapuram et al., 2025) implemented a sigmoid variant of attention in a hardware-friendly way, achieving more than 15% faster inference without sacrificing accuracy. These approaches illustrate how relatively small design changes can make long-context inference practical.
A complementary direction is cache optimization. DuoAttention (Xiao et al., 2025) splits heads into retrieval and streaming roles, keeping full key–value caches only for retrieval to reduce memory. RetroAttention (Choi et al., 2025) refreshes caches with new entries to improve coverage of early tokens. FlexPrefill (Lai et al., 2025) adapts the prefill stage by selecting query-aware sparse indices, balancing accuracy and efficiency during initialization. Together, these methods show how careful memory management can extend usable context length without prohibitive compute costs.
@@ C.1 Datasets
All experiments use subsets of FineWeb-Edu (Lozhkov et al., 2024). We train on 10B-token (10BT) and 100B-token (100BT) pools; preprocessing applies a global shuffle (seed=42). The 15B and 30B budgets are obtained deterministically by taking the first 15B / 30B tokens from the start of the shuffled 100BT subset. Ablation runs use the full 10BT pool, which is shuffled once in preprocessing and then used with a fixed order during training. For all model sizes and runs the training data and its order are identical (the same shuffled are used), ensuring data-level consistency across experiments.
@@ C.2 Implementation details
All models and baselines are implemented on top of the FLA (Yang and Zhang, 2024) and Flame (Zhang and Yang, 2025) stack, which provides Triton-based efficient-attention kernels and standard LM training utilities. For sigmoid-attention experiments we replace the FlashAttention kernel with the authors’ Flash-Sigmoid implementation integrated into the same stack.
@@ C.4 Benchmarks
Wikitext (Merity et al., 2017): A large linguistic corpus extracted from Wikipedia articles, containing over 100 million word tokens. It tests a model’s ability to predict the next word in a passage of text.
Lambada (Paperno et al., 2016): The LAmBdA dataset tests a model’s capability of using broad discourse context to predict the last word of a passage extracted from books. It contains over 60,000 examples.
PIQA (Bisk et al., 2020): The Physical Interaction: Question Answering (PIQA) dataset tests commonsense reasoning about physical interactions between two entities. It contains 16,113 multiple choice questions generated from crowd-sourcing.
Hellaswag (Zellers et al., 2019): The HellaSwag dataset consists of 70,000 multiple choice questions about inferring what might happen next in a story. It requires commonsense reasoning to choose the most plausible ending.
WinoGrande (Sakaguchi et al., 2021): The WinoGrande dataset tests coreference resolution and commonsense reasoning with 44,000 examples obtained from books and websites.
ARC (Clark et al., 2018): The AI2 Reasoning Challenge (ARC) dataset contains 7,787 genuine grade-school level, multiple-choice science questions, grouped into an Easy Set (ARC-e) and a Challenge Set (ARC-c).
SIQA (Sap et al., 2019): The Social Interaction QA (SIQA) dataset contains 15,554 multiple choice questions that describe situations about people’s social interactions.
BoolQ (Clark et al., 2019): The Boolean Questions (BoolQ) dataset contains 15,942 English yes/no questions sampled from Google search queries to test a model’s ability to answer simple questions.